# From Interpolating Formulas to Separating Languages and Back Again

### v1

@@ 3 Interpolant Existence in Logics without Craig Interpolation Property
  • propositional modal logics determined by linear frames such as $\mathsf{S4.3}$ and $\mathsf{K4.3}$ [70], the logic $\mathsf{wK4}$ (weak $\mathsf{K4}$) of derivative spaces [71] and $\mathsf{Alt}_{n}$ of $n$-ary trees (for $n\geq 3$[58], whose lack of the CIP was shown in [64, 45, 114, 58];
@@ 3.1 Interpolant Existence in Normal Modal Logics
$(iii)$ Finally, suppose $\varphi=\Diamond(p\land q)\land\Diamond(p\land\neg q)$ and $\psi=\neg[\Diamond(\neg p\land r)\land\Diamond(\neg p\land\neg r)]$ with common signature $\sigma=\{p\}$. The reader can readily check that $\varphi\to\psi$ is true in all models based on trees of outdegree 3, and so $\varphi\models_{\mathsf{Alt}_{3}}\psi$. On the other hand, in the picture below, $\mathfrak{M}_{\varphi},w_{\varphi}\models\varphi$, $\mathfrak{M}_{\psi},w_{\psi}\models\neg\psi$, and the relation $\boldsymbol{\beta}$ shown by the dotted lines is a $\sigma$-bisimulation between $\mathfrak{M}_{\varphi}$ and $\mathfrak{M}_{\psi}$ with $w_{\varphi}\boldsymbol{\beta}w_{\psi}$. It follows that $\varphi$ and $\psi$ do not have an interpolant in $\mathsf{Alt}_{3}$. $\dashv$
@@ 3.2 The Finite Bisimilar Model Property of Modal Logic with Nominals
We illustrate the reason for this higher complexity of the IEP by sketching the idea of the lower bound construction. The proof is by reduction of the NExpTime-complete $2^{n}\times 2^{n}$-tiling problem to the complement of the IEP for $\mathsf{K^{nom}}$. Given any $n<\omega$ and a finite set $T$ of tile types, we define formulas $\varphi_{n,T}$ and $\psi_{n,T}$ of polynomial size in $|T|$ and $n$ such that $\varphi_{n,T}\to\psi_{n,T}\in\mathsf{K^{nom}}$ and $T$ can tile a $2^{n}\times 2^{n}$-grid iff $\varphi_{n,T}$ and $\psi_{n,T}$ do not have an interpolant in $\mathsf{K^{nom}}$. The common signature $\sigma_{n,T}$ of $\varphi_{n,T}$ and $\psi_{n,T}$ contains $O(n)$-many variables that are used to speak about binary representations of cells in the $2^{n}\times 2^{n}$ grid; $\sigma_{n,T}$ also contains variables for the tile types in $T$. The formula $\varphi_{n,T}$ contains the conjunct $\Diamond^{2n}\boldsymbol{i}\wedge\Box^{2n}\boldsymbol{i}$, which is true in a pointed model $\mathfrak{M},w$ iff the only world accessible from $w$ in $2n$ steps in $\mathfrak{M}$ is where nominal $\boldsymbol{i}$ holds. The formula $\psi_{n,T}$ contains standard conjuncts with $O(n)$-many non-$\sigma_{n,T}$ variables that generate a full binary tree of depth $2n$ [27, 19] whose leaves $u_{1},\dots,u_{2^{2n}}$ are regarded as the cells of the grid. Now, suppose that $\mathfrak{M}_{1},w_{1}\sim_{\mathsf{ML\!^{nom}}(\sigma_{n})}\mathfrak{M}_{2},w_{2}$, $\mathfrak{M}_{1},w_{1}\models\varphi_{n,T}$ and $\mathfrak{M}_{2},w_{2}\models\psi_{n,T}$. Then all the leaves $u_{l}$ of that full binary tree of depth $2n$ in $\mathfrak{M}_{2}$ are $\sigma_{n,T}$-bisimilar to the world where $\boldsymbol{i}$ is true in $\mathfrak{M}_{1}$, and so all trees with the roots $u_{1},\dots,u_{2^{2n}}$ in $\mathfrak{M}_{2}$ are $\sigma_{n,T}$-bisimilar to each other. The other conjuncts of $\psi_{n,T}$ make sure that each tree rooted at $u_{l}$ describes the tile types of the grid cell represented by $u_{l}$ and of its neighbours in the grid; the $\sigma_{n,T}$-bisimilarity of these trees ensures that these local descriptions are consistent.
@@ 4 Interpolation, Separation and Definability in Weaker Languages
(where $\exists^{=1}y\,\alpha(y)$ says that there exists exactly one $y$ for which $\alpha(y)$ holds trues). Clearly, $\varphi\models\neg\psi$ and the $\mathsf{ML}$-formula $\chi=\Box A\vee\Box\neg A$ separates $\varphi(x)$ and $\psi(x)$. For $\psi^{\prime}(x)=\exists^{=2}y\,R(x,y)$, we also have $\varphi\models\neg\psi^{\prime}$, but $\varphi(x)$ and $\psi^{\prime}(x)$ are not $\mathsf{ML}$-separable as follows from Theorem 4.21 below. $\dashv$
@@
are $\mathsf{ML}$-definable. Indeed, let $T$ be the set of all conjunctions $B_{1}\wedge\dots\wedge B_{n}$, where $B_{i}\in\{A_{i},\neg A_{i}\}$, for $i=1,\dots,n$. Then each $\varphi_{n}$ is equivalent to $\chi_{n}=\bigvee_{t\in T}(t\wedge\Diamond t)$. It might be of interest to note that these $\chi_{n}$ are of exponential-size in $n$ and there are no $\mathsf{ML}$-formulas $\chi_{n}$ that are equivalent to $\varphi_{n}$ and have polynomial-size $|\textit{sub}(\chi_{n})|$ [68].
@@ 4.1 Definability and Separation in Fragments of First-Order Logic
  • $\varphi(x)$ is invariant under $\mathsf{ML}(\sigma)$-bisimulations, for $\sigma=\textit{sig}(\varphi)$, which means that, for any pointed models $\mathfrak{M}_{i},u_{i}$, $i=1,2$, whenever $\mathfrak{M}_{1},u_{1}\sim_{\mathsf{ML}(\sigma)}\mathfrak{M}_{2},u_{2}$, then $\mathfrak{M}_{1}\models\varphi(u_{1})$ iff $\mathfrak{M}_{2}\models\varphi(u_{2})$.
@@
While definability has been studied in depth for many pairs $L/L^{\prime}$ of logics, the work on $L/L^{\prime}$-separation has only just begun. Computationally, separation is often much harder than definability. For example, using the fact that $\mathsf{ML}$-indistinguishability in Theorem 4.21 can be replaced by $\mathsf{ML}$-bisimilarity, one can show that $\mathsf{C^{2}}/\mathsf{ML}$-separation is undecidable by a reduction of Minsky machines. It follows that Theorem 4.26 does not generalise to separation. For fragments of $\mathsf{C^{2}}$, however, a few decidability results are known. For instance, by proving the FBMP for $\mathsf{GML}$-formulas and $\mathsf{ML}$-bisimulations using an approach similar to the proof of Proposition 3.11, one can show that $\mathsf{GML}/\mathsf{ML}$-separation is NExpTime-complete—more complex than $\mathsf{GML}/\mathsf{ML}$-definability but still decidable.
@@ 4.2 Definability and Separation for Logics with Second-Order Quantifiers and Fixed-Point Operators
In fact, $\mu\mathsf{ML}$ is powerful enough to express the operators of temporal logics $\mathsf{LTL}$ (see Section 6 below), $\mathsf{CTL}$, $\mathsf{CTL}^{*}$ and propositional dynamic logic $\mathsf{PDL}$ (in which case a multimodal $\mu\mathsf{ML}$ is required).
@@
For $\varphi(P,x)=A(x)\lor\exists y\,\big{(}R(x,y)\land P(y)\big{)}$, we have $\mathfrak{A}\models\mu P.\varphi(a)$ iff $\mathfrak{A}$ contains an $R$-path from $a$ to some $b\in{\sf dom}(\mathfrak{A})$ with $\mathfrak{A}\models A(b)$. Thus, the formula $\mu P.\varphi(x)$ is equivalent to the monadic datalog query $(\Pi_{\varphi},P(x))$, in which the program $\Pi_{\varphi}$ consists of two rules: $P(x)\leftarrow A(x)$ and $P(x)\leftarrow R(x,y)\land P(y)$.
A formula $\varphi(P,x)$ is called bounded if there is some $n<\omega$ such that, for every structure $\mathfrak{A}$, the closure ordinal of $f_{\varphi}$ applied to $\mathfrak{A}$ does not exceed $n$. It is readily sees that if $\varphi(P,x)$ is bounded, then $\mu P.\varphi(P,x)$ is $\mathsf{FO}$-definable. The classical Barwise-Moschovakis theorem below states that the converse also holds for $\mathsf{FO}$:
@@
One can check that, for $\psi(P,x)$ depicted below, the closure ordinal of $f_{\psi}$ is 2. The
@@
formula $\mu P.\psi(P,x)$ corresponds to the datalog query $(\Pi_{\psi},P(x))$, where $\Pi_{\psi}$ contains the rules
@@
and so this datalog query can be rewritten to an equivalent $\mathsf{FO}$-query.
@@ 5 Definability and Separation in Formal Languages
Automata-theoretic approach to regular languages. It is well known that a language is regular iff it can be specified by a deterministic finite automaton (DFA) or, equivalently, by a nondeterministic finite automaton (NFA). Given a DFA $\mathfrak{A}$, every word $w\in A^{+}$ determines a state-to-state transition function $\delta_{w}$ in $\mathfrak{A}$: for any state $q$ in the DFA, $\delta_{w}(q)$ is the state $\mathfrak{A}$ reaches after reading $w$ when started at $q$. We say that $\delta_{w}$ has a non-trivial cycle if there exist a state $q$ and $n$, $1<n<\omega$, such that $\delta_{w}(q)\neq q$ and $\delta_{w^{n}}(q)=q$. A DFA is called counter-free if no $\delta_{w}$ has a non-trivial cycle, for any $w\in A^{+}$.
@@
We say that a language $\boldsymbol{L}$} is recognised by a semigroup $\boldsymbol{S}$} if there exist a semigroup homomorphism $\alpha\colon A^{+}\to{\boldsymbol{S}}$ and some $F\subseteq S$ such that ${\boldsymbol{L}}=\alpha^{-1}(F)$. It is easy to see that if $\boldsymbol{L}$} is recognised by a finite semigroup, then $\boldsymbol{L}$} is regular. On the other hand, from any NFA specifying a language $\boldsymbol{L}$}, one can obtain a ‘canonical’ finite semigroup recognising $\boldsymbol{L}$}, called the syntactic semigroup of $\boldsymbol{L}$}. So, we obtain:
@@
Languages over infinite words. Given an alphabet $A$, we denote by $A^{\omega}$ the set of all infinite words, called $\omega$-words, over $A$. A subset ${\boldsymbol{L}}\subseteq A^{\omega}$ is called an $\omega$-language over $A$. The notions of regularity and star-freeness of $\omega$-languages can be defined by appropriately adjusting the closure conditions given above for finite-word languages. The automata-theoretic, algebraic and logical approaches can be adapted as well. Regular $\omega$-languages are those that are specified by nondeterministic Büchi automata (NBAs[25]. Theorem 5.37 also holds in this setting: regular $\omega$-languages coincide with those that are recognised by finite $\omega$-semigroups (an appropriate generalisation of finite semigroups) [87]. Also, any $\omega$-word $w\in A^{\omega}$ can be regarded as an $\mathsf{FO}$-structure in the signature $\sigma_{A}$ whose underlying strict linear order is isomorphic to $(\omega,<)$, and Theorem 5.38 also holds: An $\omega$-language over $A$ is regular iff it is definable over $(\omega,<)$ by an $\mathsf{MSO}$-sentence in the signature $\sigma_{A}$ [25].
@@ 5.1 Definability
Languages over infinite words. $\mathsf{FO}(<)$-definable $\omega$-languages also have equivalent characterisations that are similar to those in Theorem 5.39. The equivalence of $\mathsf{FO}(<)$-definability and star-freeness is shown in [72, 106]. Given any NBA specifying an $\omega$-language $\boldsymbol{L}$}, one can obtain a ‘canonical’ structure (the syntactic $\omega$-semigroup of $\boldsymbol{L}$}[111], whose aperiodicity (that is, a property similar to aperiodicity of semigroups) is equivalent to $\boldsymbol{L}$} being star-free [86]. The equivalence of $\mathsf{FO}(<)$-definability and being specified by deterministic counter-free Büchi automata with Rabin–Streett acceptance conditions is shown in [107], based on earlier work in [80]. The survey [37] also covers several other equivalent characterisations.
@@ 5.2 Separation
Deciding $\mathsf{FO}(<)$-separability of two regular languages is in ExpTime in the size of their recognising semigroups, and so in $2\textsc{ExpTime}$ in the size of the DFAs specifying the given languages.
@@
The recognising homomorphism $\alpha$ maps any $w\in A^{+}$ to the state-to-state transition function $\delta_{w}$. It is easy to see that $S=\{\delta_{a},\delta_{b},\delta_{aa},\delta_{ab},\delta_{ba},\delta_{aba},\delta_{bab},\delta_{abab},\delta_{baba}\},$ with idempotent elements $\delta_{abab}$, $\delta_{baba}$ and $\delta_{aa}$ (with $\delta_{aa}$ being the constant $\bot$-function), and $i_{{\boldsymbol{S}}}=2$. Further, ${\boldsymbol{L}}_{1}=\alpha^{-1}(\{\delta_{abab}\})$ and ${\boldsymbol{L}}_{2}=\alpha^{-1}(\{\delta_{baba}\})$. Also, it is not hard to see that the only non-singleton elements of $\boldsymbol{S}$$^{\dagger}$ are $\{\delta_{abab},\delta_{ab}\}$, $\{\delta_{a},\delta_{aba}\}$, $\{\delta_{baba},\delta_{ba}\}$, and $\{\delta_{b},\delta_{bab}\}$. Therefore, $\{\delta_{abab},\delta_{baba}\}\notin{\boldsymbol{S}}^{\dagger}$, and so ${\boldsymbol{L}}_{1}$ and ${\boldsymbol{L}}_{2}$ are $\mathsf{FO}(<)$-separable by the criterion (1).
@@
The recognising homomorphism $\alpha$ maps any $w\in A^{+}$ to the state-to-state transition function $\delta_{w}$. It is easy to see that ${\boldsymbol{L}}_{1}=\alpha^{-1}(\{\delta_{b^{2}a}\})$ and ${\boldsymbol{L}}_{2}=\alpha^{-1}(\{\delta_{b},\delta_{b^{2}ab}\})$. We claim that $\{\delta_{b^{2}a},\delta_{b^{2}ab}\}$ in $\boldsymbol{S}$$^{\dagger}$, and so ${\boldsymbol{L}}_{1}$ and ${\boldsymbol{L}}_{2}$ are not $\mathsf{FO}(<)$-separable by the criterion (1). Indeed, observe that $i_{{\boldsymbol{S}}}=3$, so $\{\delta_{a}\}^{i_{{\boldsymbol{S}}}}=\{\delta_{a}\}$ and $\{\delta_{a}\}^{i_{{\boldsymbol{S}}}+1}=\delta_{a^{2}}$. Thus, $\{\delta_{a},\delta_{a^{2}}\}\in{\boldsymbol{S}}^{\dagger}$ by C3. Then, by C2, we obtain that $\{\delta_{a},\delta_{a^{2}}\}\mathop{\cdot}\{\delta_{b}\}=\{\delta_{ab},\delta_{a^{2}b}\}\in{\boldsymbol{S}}^{\dagger}$. Let $X=\{\delta_{ab},\delta_{a^{2}b}\}$. As $\delta_{a^{2}b}\mathop{\cdot}\delta_{ab}\mathop{\cdot}\delta_{a^{2}b}=\delta_{bab^{2}}$ and $\delta_{a^{2}b}\mathop{\cdot}\delta_{ab}\mathop{\cdot}\delta_{a^{2}b}\mathop{\cdot}\delta_{ab}=\delta_{bab}$, we have $\delta_{bab^{2}},\delta_{bab}\in X^{i_{{\boldsymbol{S}}}}\cup X^{i_{{\boldsymbol{S}}}+1}$, and so $\{\delta_{bab^{2}},\delta_{bab}\}\in{\boldsymbol{S}}^{\dagger}$ by C3 and C1. And as $\delta_{b}\mathop{\cdot}\delta_{bab^{2}}\mathop{\cdot}\delta_{a}=\delta_{b^{2}a}$ and $\delta_{b}\mathop{\cdot}\delta_{bab}\mathop{\cdot}\delta_{aa}=\delta_{b^{2}ab}$, we obtain that $\{\delta_{b^{2}a},\delta_{b^{2}ab}\}\in{\boldsymbol{S}}^{\dagger}$ by C2 and C1, as required.
@@
Languages over infinite words. According to [91], Theorems 5.41 and 5.45 also hold for $\omega$-languages.

### v2

@@ 3 Interpolant Existence in Logics without Craig Interpolation Property
  • propositional modal logics determined by linear frames such as $\mathsf{S4.3}$ and $\mathsf{K4.3}$ [70], the logic $\mathsf{wK4}$ (weak $\mathsf{K4}$) of derivative spaces [71] and $\mathsf{Alt}_{n}$ of $n$-ary trees (for $n\geq 3$) [58], whose lack of the CIP was shown in [64, 45, 114, 58];
@@ 3.1 Interpolant Existence in Normal Modal Logics
[…]
@@ 3.2 The Finite Bisimilar Model Property of Modal Logic with Nominals
We illustrate the reason why the complexity of the IEP for $\mathsf{K^{nom}}$ is higher than the complexity of its decision problem by sketching the idea of the lower bound construction. The proof is by reduction of the NExpTime-complete $2^{n}\times 2^{n}$-tiling problem to the complement of the IEP for $\mathsf{K^{nom}}$. Given any $n<\omega$ and a finite set $T$ of tile types, we define formulas $\varphi_{n,T}$ and $\psi_{n,T}$ of polynomial size in $|T|$ and $n$ such that $\varphi_{n,T}\to\psi_{n,T}\in\mathsf{K^{nom}}$ and $T$ can tile a $2^{n}\times 2^{n}$-grid iff $\varphi_{n,T}$ and $\psi_{n,T}$ do not have an interpolant in $\mathsf{K^{nom}}$. The common signature $\sigma_{n,T}$ of $\varphi_{n,T}$ and $\psi_{n,T}$ contains $O(n)$-many variables that are used to speak about binary representations of cells in the $2^{n}\times 2^{n}$ grid; $\sigma_{n,T}$ also contains variables for the tile types in $T$. The formula $\varphi_{n,T}$ contains the conjunct $\Diamond^{2n}\boldsymbol{i}\wedge\Box^{2n}\boldsymbol{i}$, which is true in a pointed model $\mathfrak{M},w$ iff the only world accessible from $w$ in $2n$ steps in $\mathfrak{M}$ is where nominal $\boldsymbol{i}$ holds. The formula $\psi_{n,T}$ contains standard conjuncts with $O(n)$-many non-$\sigma_{n,T}$ variables that generate a full binary tree of depth $2n$ [27, 19] whose leaves $u_{1},\dots,u_{2^{2n}}$ are regarded as the cells of the grid. Now, suppose that $\mathfrak{M}_{1},w_{1}\sim_{\mathsf{ML\!^{nom}}(\sigma_{n})}\mathfrak{M}_{2},w_{2}$, $\mathfrak{M}_{1},w_{1}\models\varphi_{n,T}$ and $\mathfrak{M}_{2},w_{2}\models\psi_{n,T}$. Then all the leaves $u_{l}$ of that full binary tree of depth $2n$ in $\mathfrak{M}_{2}$ are $\sigma_{n,T}$-bisimilar to the world where $\boldsymbol{i}$ is true in $\mathfrak{M}_{1}$, and so all trees with the roots $u_{1},\dots,u_{2^{2n}}$ in $\mathfrak{M}_{2}$ are $\sigma_{n,T}$-bisimilar to each other. The other conjuncts of $\psi_{n,T}$ make sure that each tree rooted at $u_{l}$ describes the tile types of the grid cell represented by $u_{l}$ and of its neighbours in the grid; the $\sigma_{n,T}$-bisimilarity of these trees ensures that these local descriptions are consistent.
@@ 4 Interpolation, Separation and Definability in Weaker Languages
[…]
@@
[…]
@@ 4.1 Definability and Separation in Fragments of First-Order Logic
  • $\varphi(x)$ is invariant under $\sigma$-bisimulations, for $\sigma=\textit{sig}(\varphi)$, which means that, for any pointed models $\mathfrak{M}_{i},u_{i}$, $i=1,2$, whenever $\mathfrak{M}_{1},u_{1}\sim_{\mathsf{ML}(\sigma)}\mathfrak{M}_{2},u_{2}$, then $\mathfrak{M}_{1}\models\varphi(u_{1})$ iff $\mathfrak{M}_{2}\models\varphi(u_{2})$.
@@
While definability has been studied in depth for many pairs $L/L^{\prime}$ of logics, the work on $L/L^{\prime}$-separation has only just begun. Computationally, separation is often much harder than definability. For example, using the fact that $\mathsf{ML}$-indistinguishability in Theorem 4.21 can be replaced by $\mathsf{ML}$-bisimilarity, one can show that $\mathsf{C^{2}}/\mathsf{ML}$-separation is undecidable by a reduction of Minsky machines. It follows that Theorem 4.26 does not generalise to separation. For fragments of $\mathsf{C^{2}}$, however, a few decidability results are known. For instance, by proving the FBMP for $\mathsf{GML}$-formulas and $\mathsf{ML}$-bisimulations using an approach similar to the proof of Proposition 3.11, one can show that $\mathsf{GML}/\mathsf{ML}$-separation is coNExpTime-complete—more complex than $\mathsf{GML}/\mathsf{ML}$-definability but still decidable.
@@ 4.2 Definability and Separation for Logics with Second-Order Quantifiers and Fixed-Point Operators
[…]
@@
[…]
A formula $\varphi(P,x)$ is called bounded if there is some $n<\omega$ such that, for every structure $\mathfrak{A}$, the closure ordinal of $f_{\varphi}$ applied to $\mathfrak{A}$ does not exceed $n$. It is readily seen that if $\varphi(P,x)$ is bounded, then $\mu P.\varphi(P,x)$ is $\mathsf{FO}$-definable. The classical Barwise-Moschovakis theorem below states that the converse also holds for $\mathsf{FO}$:
@@
One can check that, for $\psi(P,x)$ depicted below, the closure ordinal of $f_{\psi}$ is 2:
@@
The formula $\mu P.\psi(P,x)$ corresponds to the datalog query $(\Pi_{\psi},P(x))$, where $\Pi_{\psi}$ contains the rules
@@
[…]
@@ 5 Definability and Separation in Formal Languages
Automata-theoretic approach to regular languages. It is well known that a language is regular iff it can be specified by a deterministic finite automaton (DFA) or, equivalently, by a nondeterministic finite automaton (NFA). Given an NFA $\mathfrak{A}$, every word $w\in A^{+}$ determines a binary transition relation $\Delta_{w}$ on the states of $\mathfrak{A}$: it consists of all those state-pairs $(q,q^{\prime})$, for which $\mathfrak{A}$ can reach $q^{\prime}$ having started at $q$ and read $w$. If $\mathfrak{A}$ is a DFA, then each $\Delta_{w}$ is actually a state-to-state transition function $\delta_{w}$. We say that such a $\delta_{w}$ has a non-trivial cycle if there exist a state $q$ and $n$, $1<n<\omega$, such that $\delta_{w}(q)\neq q$ and $\delta_{w^{n}}(q)=q$. A DFA is called counter-free if no $\delta_{w}$ has a non-trivial cycle, for any $w\in A^{+}$.
@@
We say that a language $\boldsymbol{L}$} is recognised by a semigroup $\boldsymbol{S}$} if there exist a semigroup homomorphism $\alpha\colon A^{+}\to{\boldsymbol{S}}$ and some $F\subseteq S$ such that ${\boldsymbol{L}}=\alpha^{-1}(F)$. It is easy to see that if $\boldsymbol{L}$} is recognised by a finite semigroup, then $\boldsymbol{L}$} is regular. On the other hand, any language $\boldsymbol{L}$} that is specified by an NFA $\mathfrak{A}$ is always recognised by the transition semigroup of $\mathfrak{A}$. As the latter is finite (of size at most $2^{n^{2}}$, for the number $n$ of states in $\mathfrak{A}$), we obtain:
@@
In fact, for any regular language $\boldsymbol{L}$}, one can obtain a ‘canonical’ finite semigroup recognising $\boldsymbol{L}$}, called the syntactic semigroup of $\boldsymbol{L}$}: it is the transition semigroup of the minimal DFA specifying $\boldsymbol{L}$}.
@@
Languages over infinite words. Given an alphabet $A$, we denote by $A^{\omega}$ the set of all infinite words, called $\omega$-words, over $A$. A subset ${\boldsymbol{L}}\subseteq A^{\omega}$ is called an $\omega$-language over $A$. The notions of regularity and star-freeness of $\omega$-languages can be defined by appropriately adjusting the closure conditions given above for finite-word languages. The automata-theoretic, algebraic and logical approaches can be adapted as well. Regular $\omega$-languages are those that are specified by nondeterministic Büchi automata (NBAs) [25]. Theorem 5.37 also holds in this setting: regular $\omega$-languages coincide with those that are recognised by finite $\omega$-semigroups (an appropriate generalisation of finite semigroups) [87]. Also, any $\omega$-word $w\in A^{\omega}$ can be regarded as an $\mathsf{FO}$-structure in the signature $\sigma_{A}$ whose underlying strict linear order is isomorphic to $(\omega,<)$, and Theorem 5.38 also holds: An $\omega$-language over $A$ is regular iff it is definable over $(\omega,<)$ by an $\mathsf{MSO}$-sentence in the signature $\sigma_{A}$ [25].
@@ 5.1 Definability
Languages over infinite words. $\mathsf{FO}(<)$-definable $\omega$-languages also have equivalent characterisations that are similar to those in Theorem 5.39. The equivalence of $\mathsf{FO}(<)$-definability and star-freeness is shown in [72, 106]. Given any NBA specifying an $\omega$-language $\boldsymbol{L}$}, one can obtain a ‘canonical’ structure (the syntactic $\omega$-semigroup of $\boldsymbol{L}$}) [111], whose aperiodicity (that is, a property similar to aperiodicity of semigroups) is equivalent to $\boldsymbol{L}$} being star-free [86]. The equivalence of $\mathsf{FO}(<)$-definability and being specified by deterministic counter-free Büchi automata with Rabin–Streett acceptance conditions is shown in [107], based on earlier work in [80]. The survey [37] also covers several other equivalent characterisations.
@@ 5.2 Separation
Deciding $\mathsf{FO}(<)$-separability of two regular languages is in ExpTime in the size of their recognising semigroups, and so in $2\textsc{ExpTime}$ in the size of the NFAs specifying the given languages.
@@
[…]
@@
[…]
@@
Languages over infinite words. According to [91], the analogues of Theorems 5.41, 5.42, and 5.45 also hold for $\omega$-languages.